Please help.
2) Use the geometric sequence of numbers 1, 3, 9, 27, … to find the following:
a) What is r, the ratio between 2 consecutive terms?
Show work in this space.



b) Using the formula for the nth term of a geometric sequence, what is the 10th term?
Answer: the sum of the first 10 terms is: Show work in this space.

c) Using the formula for the sum of a geometric series, what is the sum of the first 10 terms?
Answer: Show work in this space.

Answer by venugopalramana(3286)

Please help.
2) Use the geometric sequence of numbers 1, 3, 9, 27, … to find the following:
a) What is r, the ratio between 2 consecutive terms?
Show work in this space.
3/1=9/3=27/9=3=R=COMMON RATIO


b) Using the formula for the nth term of a geometric sequence, what is the 10th term?
TN=A*R^(N-1)
T10=1*3^(10-1)=3^9
Answer: the sum of the first 10 terms is: Show work in this space.
SN=A*{(R^N -1)/(R-1)}
S10=1*{(3^10 -1)/(3-1)}=(3^10 -1)/2
c) Using the formula for the sum of a geometric series, what is the sum of the first 10 terms?
Answer: Show work in this space.